A projectile fired at O follows a parabolic trajectory, given in parametric form by:

x = 86t and y = 96t – 4.9t2

where x and y are measured in meters and t in seconds. Determine:

(a) ax and ay at any time t

(b) vx and vy at any time t

(c) initial velocity, vo (velocity at pt O)

(d) velocity when t = 5s (hint: get first vx and vy at t = 5s)

(e) maximum height h

(f) horizontal distance L
1
Expert's answer
2020-04-27T10:00:08-0400

Solution.

"x = 86t;" "y = 96t-4.9t^2;"

Since the trajectory of motion is a parabola, the motion of the projectile is described by the formulas:

"x = \upsilon_0"x"t;" "y = \upsilon_0"y "t + a_yt^2\/2;"

a) "a_x = 0; a_y = -9.8m\/s^2;"

b) "\upsilon_x = 86m/s ;""\upsilon_y = 96-9.8t;"


c) "\upsilon_0 =\sqrt{\upsilon_x^2 + \upsilon_y^2 };" "\upsilon_0 = \sqrt{86^2 + 96^2 } = 128.9 m/s;"

d) "\upsilon_x = 86 m/s; \upsilon_y = 96 - 9.8\sdot 5 = 47m/s;"

e) "h_m = -\upsilon_0"y2 /2"\sdot a_y;"

"h_m = -96^2/2\sdot(-9.8) = 470.2m;"

f) "L =\u03c5_0\n\u200b"x"t_m;" "t_m =2\sdot(-\upsilon"0y"\/a_y);" "t_m = 2\sdot(-96/-9.8) =19.6s;"

"L = 86\sdot19.6 = 1684.9m;"

Answer: a) "a_x = 0; a_y = -9.8 m\/s^2;"

b) "\upsilon_x = 86m/s ;" "\upsilon_y = 96-9.8t;"

c) "\upsilon_0 = 128.9m/s;"

d) "\upsilon_x = 86 m/s; \upsilon_y = 47m/s;"

e)"h_m = 470.2m;"

f)"L = 1684.9m;"